# Calibration Method of Spacecraft–Inertial Sensor Center-of-Mass Offset for the Taiji Gravitational Wave Detection Mission under Science Mode

### v1

@@ Abstract
Accurately calibrating the center-of-mass (CoM) offset between the spacecraft (SC) and the inertial sensor test mass (TM) is crucial for space-based gravitational-wave (GW) antennas, such as LISA and Taiji. Current calibration methods require additional spacecraft maneuvers that disrupt science data continuity and inter-satellite links, compromising the coherence of gravitational wave signals. Here, we present a maneuver-free calibration scheme that directly estimates the CoM offset vector using only standard science-mode measurements from inertial sensors, interferometers, and differential wavefront sensors. By embedding the CoM offset induced coupling acceleration as an extended state in a model-based adaptive Kalman filter, we achieve estimation accuracy of 0.01–1.5 mm across all axes with a maximum error below 1%. This approach enables continuous, high-precision calibration during nominal observation runs, ensuring continuous and coherent gravitational wave data collection while maintaining the required precision, and also facilitating advanced DFACS functions such as performance evaluations and fault diagnosis. For LISA-like missions, where data continuity is paramount for detecting faint gravitational wave signals, this method will enhance scientific output and reliability.
@@ I Introduction
The Laser Interferometer Space Antenna (LISA) mission [LISA2024] was officially approved by the European Space Agency in January 2024. As the most developed space mission, LISA had motivated various space gravitational wave (GW) detection projects since the 2000s, including DECIGO [DECIGOsato2017status, DECIGOando2010decigo, DECIGOkawamura2011japanese], Taiji [TAIJI_OVERVIEW1, hu2017taiji], TainQin [TQ_OVERVIEW, TQ_liu2020science, TQ_mei2021tianqin], ASTROD [ni2002astrod, ni2008astrod] and etc. Compared with the ground-based large interferometers, that the LIGO-Virgo-KAGRA collaboration [LIGO2016, caron1997virgo, kagra_sekiguchi2012current, lisa_virgo_kagra_abbott2020prospects], space-borne antennas will be free from Newtonian gradient noises situated near the 10 Hz band and limitations on the baselines. The LISA and the mentioned LISA-like missions will have sensitive frequency bands from 0.1 mHz to 1 Hz and will encompass scientific objectives related to much heavier sources such as coalescing super-massive black holes and extreme-mass-ratio inspirals, and also stochastic GW background from the early Universe.
@@
“The Taiji program in space” is a space-borne GW detection mission of China, released by the Chinese Academy of Sciences in 2016. Following its roadmap [luo2020taiji], the first technology demonstration satellite, that the Taiji-1, was launched in August 2019 and has successfully validated the related key technologies including the high precision inertial sensor (IS, also called GRS, Gravitational Reference Sensor) [GRS_2022], laser interferometer in space, drag-free controls and ultra-clean and stable satellite platform [TAIJI-1_OVERVIEW]. The Taiji mission, as one of the representative LISA-like missions, will launch three spacecraft (SC) to heliocentric orbits, forming a nearly equilateral triangular constellation with armlength $\sim 3\times 10^{6}$ km. It is expected that the science operations of LISA and Taiji may overlap in the 2030s and studies of the joint data analysis has already aroused more interest [lisataiji_wang2021alternative, lisataiji_cornish_shuman2022massive, LISATAIJI_ruan2020lisa, LISATAIJI_orlando2021measuring].
@@
High precision IS are of key payloads for the LISA and Taiji missions. The suspended test mass (TM) of each IS serves as the free-falling reference and also as the end mirror for the inter-satellite interferometers. The relative motions between the TMs induced by incident GWs could be measured precisely by means of the so called time-delay interferometry [TDI_tinto2021time] and therefore give rise to the detection of GW signals. The performance of the IS, or the residual acceleration noises of the TMs relative to the local inertial frame, will directly determine the sensitivity level of the space GW antennas. For the Taiji mission, according to its designed sensitivity and noise budgets, the residual acceleration noise for each TM along the sensitive axis (along with the inter-satellite laser link) of the IS should be $\leq 3\times 10^{-15}\ \mathrm{m/s^{2}/Hz^{\frac{1}{2}}}$ [luo2020brief]. Such stringent requirement will be upheld and guaranteed by the Drag-Free and Attitude Control System (DFACS) [bender2000lisa], which monitors in real-time and adjusts actively the relative displacements and attitudes between the SC and the TMs. The DFACS maintains the two TMs on-board each SC in nearly perfect free fall along their sensitive axes, while forcing them to follow the SC for the remaining degrees of freedom (including attitudes) [gath2004drag]. Given the careful control of each SC’s attitude and the breathing angle of the movable optical sub-assemblies to ensure stable inter-satellite measurement links, the attitude of the TMs will be maintained to align with the laser links, effectively acting as the free-falling end mirrors of the inter-satellite interferometer.
@@
Extensive studies on the DFACS for LISA-like missions have established a comprehensive model of the relative dynamics of the TMs and SC [gath2007drag]. Within this model, the relative center-of-mass (CoM) offset between the SC and TM (called the SC–TM CoM offset) constitutes the most significant nonlinear effect and gives rise to an important disturbance to the DFACS. The CoM offset vector and its variations will couple with the angular acceleration and angular velocity of the SC respectively, thereby introducing inertial acceleration noises in the local frame [vidano2020lisa]. Such CoM coupled noise will also introduce interference into the performance evaluation of the on-board ISs. For example, in the data analysis work of the LISA Pathfinder, the CoM offset was precisely calibrated, and its influence, that the inertial force noises, was accurately removed in obtaining the expected noise curve [LPF2018calibrating].
@@
For satellite gravity missions, such as CHAMP [reigber2002champ], GRACE [GRACE_weil1997gravity], and GFO [gfo_kornfeld2019grace], the onboard IS operated in accelerometer mode, and their readouts provided the necessary measurements of non-gravitational perturbations acting on the corresponding SCs. The CoM coupled effects will produce inertial acceleration noises directly to the measured data, degrading the scientific outputs of satellite missions with accelerometers for measuring non-gravitational orbit perturbations. Therefore, the in situ and accurate calibration of the SC–TM CoM offset is of importance [COM_CALI_TH] for those missions including GW antennas in space. Take GRACE as a representative example, one performed the CoM calibration by commanding large spacecraft attitude maneuvers with magnetorquers. The attitude jitters and the CoM coupled acceleration signals were measured with statrackers and the accelerometer, and then the offset vector was estimated using the least-squares method [wang2003study, COM_CALI_WANG_JOURNAL]. Taiji-1  [taiji2021china, COM_CALI_ZHANG, COM_CALI_WEI] and TianQin-1 (the technology demonstration satellite of TianQin project) [tq1_zhou2022non, TQ1_RESULT] adopted the similar maneuver strategy and employed Kalman filtering algorithms to estimate the CoM offset. The feasibility of applying this method to the Taiji mission was studied and verified with numerical simulations [liu2025design]. Furthermore, the TianQin team also analyzed the maneuver design criteria for the possible CoM calibration, which could maintain the inter-satellite link at the same time [wang2025research]. In short, during their commissioning and calibration phases, the maneuver-based SC–TM CoM offset calibration methods could meet the requirements of the LISA-like missions [lisa2011lisa, liu2025design].
@@
During the years-long scientific lifetime of LISA and Taiji, the SC-TM CoM offset vector could change slowly with time, which may be caused by the continuous consumption of the propellant, the structural deformation due to material aging, or the possible drift of the balance point of the TMs. Therefore, the scheduled CoM calibration in the science operation phase will be necessary in ensuring the quality of scientific data and the accurate assessment of the performance of IS [lupi2019precise]. However, the maneuver-based calibration will cost extra price when the space-borne GW antenna is performing science measurements. The rather large attitude swings (such as the 100 $\mathrm{\mu N\cdot m}$ maneuver proposed in [wang2025research]) would consume a substantial portion of the DFACS control margin and severely compromise the stability of the inter-spacecraft laser links. This contradicts the very purpose of using the H control approach in DFACS to ensure system robustness. More critically, the maneuvers and the resulting coupled acceleration noises would disrupt the continuity of the science data and break the coherence of the GW signals, thereby would increase the difficulty of source parameter estimations and may even affect the final scientific output of the missions.
@@
To meet the CoM offset calibration requirements of the Taiji mission, this paper proposes a maneuver-free scheme applicable during the science run without any disruption to the measurement system. As discussed earlier, the CoM offset induces disturbances that can substantially influence the performance of the DFACS system, which, correspondingly, will be manifested in the readout signals of the DFACS measurement subsystems—namely, the local TM laser interferometer (IFO) [OB_brzozowski2022lisa], differential wavefront sensor (DWS) [2022lisa_OB], and IS. These readouts and their correlations make it possible to identify the SC–TM CoM offset vector given the SC–TM dynamical model. We developed a high-fidelity DFACS numerical simulations and generate the synthetic readout data set. By embedding the CoM offset as an extended state, alongside the bias terms and the dominant colored noises, we reformulate this problem as an extended estimation problem that could be solved via an improved Sage–Husa adaptive Kalman filter. The proposed method enables high-precision CoM offset identification directly from unperturbed science-mode data, without requiring auxiliary maneuvers.
@@
This paper expands as follows. Section II introduces the purpose of the DFACS system, describes the physical configuration and operating principle of the IS, and presents the equations governing the SC–TM relative dynamics and spacecraft attitude motion. Section III designs and constructs a DFACS for Taiji to achieve high-precision simulation of the relative attitude/displacement between the SC-TM and the SC attitude signal in the inertial frame under science mode. Section IV presents the CoM calibration methodology, detailing a baseline approach that relies solely on DFACS-provided SC–TM relative displacement data without spacecraft maneuvers. The method is validated using 25 days of high-fidelity DFACS simulation data, and the resulting CoM offset estimates are reported. Finally, Section V concludes the paper and discusses implications for mission operations.
@@ II Spacecraft-Test Mass Dynamics Model
Figure 1 illustrates the internal layout configuration of the Taiji spacecraft. For concreteness, SC1, the first spacecraft in the constellation, is used as a representative example. Inside SC1, two TMs undergo free fall along their respective sensing axes, $\mathbf{o}_{11}$ and $\mathbf{o}_{12}$, which are separated by an angle of $60^{\circ}$. Each TM is enclosed within an electrode cage, rigidly mounted to both the SC and its associated optical bench. The IS employs sensing electrodes in the cage to measure the TM-SC relative displacement, and actuation electrodes to apply electrostatic forces on the TM.
Owing to their shared measurement principle, LISA, Taiji, and TianQin feature highly similar internal mechanical layouts [DFACSbohan2024review]. we adopt the coordinate systems established in [vidano2020lisa], which provide consistent frames for expressing forces, torques, and kinematic relationships. The key reference frames illustrated in Figure 1 are defined as follows:
@@
$\mathbf{r}_{m_{i}O_{i}}^{O_{i}}$: The relative displacement between the i-th optical platform and the i-th TM in $\text{ORF}_{i}$.
@@
As the rigorous derivations of DFACS dynamics have been well established in previous studies (e.g., [vidano2020lisa, DFACSbohan2024review]), this paper focuses on the key dynamic relationships essential for controller design and analysis. Specifically, we present: (i) the translational and rotational dynamics of the $i$-th TM relative to the $\text{ORF}_{i}$ — namely, $\mathbf{r}_{m_{i}O_{i}}^{O_{i}}$ and $\mathbf{q}_{m_{i}O_{i}}$ — under drag-free control; and (ii) the attitude dynamics of the SC relative to the CRF, described by $\mathbf{q}_{SC}$, under attitude control.
Accurate modeling of the relative dynamics between the SC and its free-falling TMs is fundamental to the DFACS and CoM calibrations for Taiji. To formulate these dynamics, we begin with the translational dynamics of the SC center of mass in the IRF:
$\ddot{\mathbf{r}}_{SI}^{I}=\boldsymbol{a}_{g}(\mathbf{r}_{SI}^{I})+m_{S}^{-1}\mathbf{T}_{S}^{I}\left(\mathbf{F}_{T}+\mathbf{d}_{S}-\sum_{i=1}^{2}\mathbf{T}_{O_{i}}^{S}\mathbf{F}_{E_{i}}\right).$ (1)
@@
where $\mathbf{r}_{\text{sun}}$ and $\mathbf{r}_{e}$ are the position vectors of the Sun and Earth’s mass centers expressed in IRF. $\mu_{s}$ is the solar gravitational constant, $\mu_{e}$ is the Earth’s gravitational constant. Similarly, the acceleration of TM in IRF is given by
@@
where $\mathbf{a}_{g}(r_{m}^{I})$ is the gravitational acceleration experienced by TM, $\mathrm{m_{m}}$ is the mass of TM, $\mathbf{F}_{\mathrm{st}}$ represents the TM stiffness force, primarily induced by the self-gravity gradient [stiffness_2009Current]. $\mathbf{d}_{m}$ denotes other environmental noise on the TM. Finally, the relative acceleration of the i-th TM with respect to the SC can be expressed as
@@
In eq(4), $\boldsymbol{\omega}_{SI}$ represents the angular velocity of the SC relative to the IRF expressed in SRF. Unless otherwise specified, all angular velocities and angular accelerations (e.g., $\boldsymbol{\omega}_{SI}$, $\dot{\boldsymbol{\omega}}_{SI}$) are expressed in their body fixed reference frame; thus, explicit superscripts (e.g., S) are omitted for brevity throughout this work.
@@
where $\mathbf{J}_{S}$ is the moment of inertia matrix of the SC, $\mathbf{M}_{T}$ is the control torque generated by the MPS, $\mathbf{D}_{S}$ denotes the environmental disturbance torque, and $\mathbf{M}_{E_{i}}$ represents the electrostatic torque applied on the $i$-th TM. Based on Eq. (5), the angular acceleration of the SC relative to the CRF can be obtained as

### v2

@@ Abstract
Accurately calibrating the center-of-mass (CoM) offset between the spacecraft (SC) and the inertial sensor test mass (TM) is crucial for space-based gravitational wave (GW) antennas, such as LISA and Taiji. Current calibration methods require additional spacecraft maneuvers that disrupt science data continuity and inter-satellite links, compromising the coherence of GW signals. Here, we present a maneuver-free calibration scheme that directly estimates the CoM offset vector using only standard science-mode measurements from inertial sensors, interferometers, and differential wavefront sensors. By embedding the CoM offset induced acceleration as an extended state in a model-based adaptive Kalman filter, we achieve estimation accuracy of 0.01–1.5 mm across all axes with a maximum error below 1%. This approach enables continuous, high-precision calibration during nominal science operations, ensuring continuous and coherent gravitational wave data collection while maintaining the required precision, and also facilitating advanced DFACS functions such as performance evaluations and fault diagnosis. For LISA-like missions, where data continuity is paramount for detecting faint and long-standing GW signals, this method will enhance scientific output and reliability.
@@ I Introduction
The Laser Interferometer Space Antenna (LISA) mission [12] was officially approved by the European Space Agency in January 2024. As the most developed space mission, LISA had motivated various space gravitational wave (GW) detection projects since the 2000s, including DECIGO [44, 4, 24], Taiji [41, 21], TainQin [30, 28, 35], ASTROD [37, 38] and etc. Compared with the ground-based large interferometers, that the LIGO-Virgo-KAGRA collaboration [1, 11, 45, 2], space-borne antennas will be free from Newtonian gradient noises situated near the 10 Hz band and limitations on the baselines. The LISA and the mentioned LISA-like missions will have sensitive frequency bands from 0.1 mHz to 1 Hz and will encompass scientific objectives related to much heavier sources such as coalescing super-massive black holes and extreme-mass-ratio inspirals, and also stochastic GW background from the early Universe.
@@
“The Taiji program in space” is a space-borne GW detection mission of China, released by the Chinese Academy of Sciences in 2016. Following its roadmap [32], the first technology demonstration satellite, that the Taiji-1, was launched in August 2019 and has successfully validated the related key technologies including the high precision inertial sensor (IS, also called GRS, Gravitational Reference Sensor) [13], laser interferometer in space, drag-free controls and ultra-clean and stable satellite platform [10]. The Taiji mission, as one of the representative LISA-like missions, will launch three spacecraft (SC) to heliocentric orbits, forming a nearly equilateral triangular constellation with armlength $\sim 3\times 10^{6}$ km. It is expected that the science operations of LISA and Taiji may overlap in the 2030s and studies of the joint data analysis has already aroused more interest [56, 46, 42, 39].
@@
High precision IS are of key payloads for the LISA and Taiji missions. The suspended test mass (TM) of each IS serves as the free-falling reference and also as the end mirror for the inter-satellite interferometers. The relative motions between the TMs induced by incident GWs could be measured precisely by means of the so called time-delay interferometry [49] and therefore give rise to the detection of GW signals. The performance of the IS, or the residual acceleration noises of the TMs relative to the local inertial frame, will directly determine the sensitivity level of the space GW antennas. For the Taiji mission, according to its designed sensitivity and noise budgets, the residual acceleration noise for each TM along the sensitive axis (along with the inter-satellite laser link) of the IS should be $\leq 3\times 10^{-15}\ \mathrm{m/s^{2}/Hz^{\frac{1}{2}}}$ ranging from 0.1 mHz to 1 Hz. [31]. Such stringent requirement will be upheld and guaranteed by the Drag-Free and Attitude Control System (DFACS) [6], which monitors in real-time and adjusts actively the relative displacements and attitudes between the SC and the TMs. The DFACS maintains the two TMs on-board each SC in nearly perfect free fall along their sensitive axes, while forcing them to follow the SC for the remaining degrees of freedom (including attitudes) [15]. Given the careful control of each SC’s attitude and the breathing angle of the movable optical sub-assemblies to ensure stable inter-satellite measurement links, the attitude of the TMs will be maintained to align with the laser links, effectively acting as the free-falling end mirrors of the inter-satellite interferometer.
@@
Extensive studies on the DFACS for LISA-like missions have established a comprehensive model of the relative dynamics of the TMs and SC [16]. Within this model, the relative center-of-mass (CoM) offset between the SC and TM (called the SC–TM CoM offset) constitutes the most significant nonlinear effect and gives rise to an important disturbance to the DFACS. The CoM offset vector and its variations will couple with the angular acceleration and angular velocity of the SC respectively, thereby introducing inertial acceleration noises in the local frame attached to the SC. Such CoM coupled noise will also introduce interference into the performance evaluation of the on-board ISs. For example, in the data analysis work of the LISA Pathfinder, the CoM offset was precisely calibrated, and its influence, that the inertial force noises, was accurately removed in obtaining the expected noise curve [5].
@@
For satellite gravity missions, such as CHAMP [40], GRACE [59], and GFO [25], the onboard IS operated in accelerometer mode, and their readouts provided the necessary measurements of non-gravitational perturbations acting on the corresponding SCs. The CoM coupled effects will produce inertial acceleration noises directly to the measured data, degrading the scientific outputs of satellite missions with accelerometers for the measuring non-gravitational orbit perturbations. Therefore, the in situ and accurate calibration of the SC–TM CoM offset is of importance [22] for those missions also including GW antennas in space. Take GRACE as a representative example, one performed the CoM calibration by commanding large spacecraft attitude maneuvers with magnetorquers. The attitude jitters and the CoM coupled acceleration signals were measured with star-trackers and the accelerometer, and then the offset vector was estimated using the least-squares method [55, 54]. Taiji-1  [60, 63, 58] and TianQin-1 (the technology demonstration satellite of the TianQin project) [64, 29] adopted the similar maneuver strategy and employed Kalman filtering algorithms to estimate the CoM offset. The feasibility of applying this method to the Taiji mission was studied and verified with numerical simulations [27]. Furthermore, the TianQin team also analyzed the maneuver design criteria for the possible CoM calibration, which could maintain the inter-satellite link at the same time [57]. In short, during their commissioning and calibration phases, the maneuver-based SC–TM CoM offset calibration methods could meet the requirements of those LISA-like missions [48, 27].
@@
During the years-long science lifetime of Taiji and LISA, the SC-TM CoM offset vector could change slowly with time, which may be caused by the continuous consumption of the propellant, the structural deformation due to material aging, or the possible drift of the balance point of the TMs. Therefore, the scheduled CoM calibration in the science operation phase will be necessary in ensuring the quality of scientific data and the accurate assessment of the performance of the ISs themselves [33]. However, the maneuver-based calibration will cost extra price when the space-borne GW antenna is performing science measurements. The rather large attitude swings (such as the 100 $\mathrm{\mu N\cdot m}$ maneuver proposed in [57]) would consume a substantial portion of the DFACS control margin and severely compromise the stability of the inter-spacecraft laser links. This contradicts the very purpose of using the H control approach in DFACS to ensure system robustness. More critically, the maneuvers and the resulting coupled acceleration noises would disrupt the continuity of the science data and break the coherence of the GW signals, thereby would increase the difficulty of source parameter estimations and may even affect the final scientific output of the missions.
@@
To meet the CoM offset calibration requirements of the Taiji mission, this paper proposes a maneuver-free scheme applicable during the science run without any disruption to the measurement system. As discussed earlier, the CoM offset induces disturbances that can substantially influence the performance of the DFACS system, which, correspondingly, will be manifested in the readout signals of the DFACS measurement subsystems—namely, the local TM laser interferometers (IFO), differential wavefront sensors (DWS) [9], and ISs. These readouts and their correlations make it possible to identify the SC–TM CoM offset vector given the SC–TM dynamical model. We developed a high-fidelity DFACS numerical simulations and generate the synthetic readout data set. By embedding the CoM offset as an extended state, alongside the bias terms and the dominant colored noises, we reformulate this problem as an extended estimation problem that could be solved via an improved Sage–Husa adaptive Kalman filter. The proposed method enables high-precision CoM offset identification directly from undisrupted science-mode data, without requiring auxiliary maneuvers.
@@
This paper expands as follows. Section II introduces the purpose of the DFACS system, describes the physical configuration and operating principle of the IS, and presents the equations governing the SC–TM relative dynamics and spacecraft attitude motion. Section III designs a DFACS for Taiji and provides a high-precision simulation of the relative attitude/displacement between the SC-TM and the SC attitude signal in the inertial frame under science mode. Section IV presents the CoM calibration methodology, detailing a novel maneuver-free approach that relies solely on DFACS-provided SC–TM relative displacement data. The method is validated using 25 days of high-fidelity DFACS simulation data, and the resulting CoM offset estimates are reported. Finally, Section V concludes the paper and discusses implications for mission operations.
@@ II Spacecraft-Test Mass Dynamics Model
Fig. 1 illustrates the internal layout configuration of the Taiji spacecraft. For concreteness, SC1 is used as a representative example. Inside SC1, two TMs undergo free fall along their respective sensing axes, $\mathbf{o}_{11}$ and $\mathbf{o}_{12}$, which are separated by an angle of approximately $60^{\circ}$. Each TM is enclosed within an electrode cage, rigidly mounted to its associated optical bench and to the SC platform. The IS employs sensing electrodes in the cage to measure the TM-SC relative displacement and rotation, and actuation electrodes to apply electrostatic forces on the TM.
Owing to their shared measurement principle, LISA, Taiji, and TianQin feature highly similar internal mechanical layouts [7]. we adopt the coordinate systems established in [52], which provide consistent frames for expressing forces, torques, and kinematic relationships. The key reference frames illustrated in Fig. 1 are defined as follows:
@@
$\mathbf{r}_{m_{i}O_{i}}^{O_{i}}$: The relative displacement between the i-th optical bench and the i-th TM in $\text{ORF}_{i}$.
@@
As the rigorous derivations of DFACS dynamics have been well established in previous studies (e.g., [52, 7]), this paper focuses on the key dynamic relationships essential for controller design and analysis. Specifically, we present: (i) the translational and rotational dynamics of the $i$-th TM relative to the $\text{ORF}_{i}$ — namely, $\mathbf{r}_{m_{i}O_{i}}^{O_{i}}$ and $\mathbf{q}_{m_{i}O_{i}}$ — under drag-free control; and (ii) the attitude dynamics of the SC relative to the CRF, described by $\mathbf{q}_{SC}$, under attitude control.
Accurate modeling of the relative dynamics between the SC and its free-falling TMs is fundamental to the DFACS and CoM calibrations for Taiji. To formulate these dynamics, we begin with the translational dynamics of the SC’s CoM in the IRF:
$\ddot{\mathbf{r}}_{SI}^{I}=\mathbf{a}_{g}(\mathbf{r}_{SI}^{I})+m_{S}^{-1}\mathbf{T}_{S}^{I}\left(\mathbf{F}_{T}+\mathbf{d}_{S}-\sum_{i=1}^{2}\mathbf{T}_{O_{i}}^{S}\mathbf{F}_{E_{i}}\right).$ (1)
@@
where $\mathbf{r}_{\text{sun}}$ and $\mathbf{r}_{e}$ are the position vectors of the Sun and Earth’s mass centers in IRF. $\mu_{s}$ is the solar gravitational constant, $\mu_{e}$ is the Earth’s gravitational constant. Similarly, the acceleration of TM in IRF is given by
@@
where $\mathbf{a}_{g}(r_{m}^{I})$ is the gravitational acceleration experienced by TM, $\mathrm{m_{m}}$ is the mass of TM, $\mathbf{F}_{\mathrm{st}}$ represents the TM stiffness force, primarily induced by the self-gravity gradient [36]. $\mathbf{d}_{m}$ denotes other environmental noise on the TM. Finally, the relative acceleration of the i-th TM with respect to the SC can be expressed as
@@
In Eq. (4), $\boldsymbol{\omega}_{SI}$ represents the angular velocity of the SC relative to the IRF expressed in SRF. Unless otherwise specified, all angular velocities and angular accelerations (e.g., $\boldsymbol{\omega}_{SI}$, $\dot{\boldsymbol{\omega}}_{SI}$) are expressed in their body fixed reference frame; thus, explicit superscripts (e.g., S) are omitted for brevity throughout this work.
@@
where $\mathbf{J}_{S}$ is the inertia moment of the SC, $\mathbf{M}_{T}$ is the control torque generated by the MPS, $\mathbf{D}_{S}$ denotes the environmental disturbance torque, and $\mathbf{M}_{E_{i}}$ represents the electrostatic torque applied on the $i$-th TM. Based on Eq. (5), the angular acceleration of the SC relative to the CRF can be obtained as